Larkspur Incorporated purchased metal to build a new roller coaster on December 31, 2020. Larkspur provided a $800,000 down payment and agreed to pay the balance in equal instalments of $280,000 every December 31 for five years. Larkspur could have received a loan from the bank for this amount at 9% interest. Larkspur prepares financial statements in accordance with IFRS.

1.

Amount used to record the note payable

2. Prepare an instalment note repayment schedule for the term of the note.

3. Prepare the journal entries that would be recorded for the purchase and for the payments and interest on December 31, 2020, 2021, 2022, 2023, 2024, and 2025
Answer #1

1)Present value of annual installment =PVA9%,5 * Annual installment

                                    = 3.88965 *280000

                                    = $ 1089102

**Find present value annuity due factor from table at 9% for 5 periods or using the formula [1/(1+i) + 1/(1+i)2 +1/(1+i)^3+ 1/(1+i) 4 +1/(1+i)^5]

2)

SCHEDULE OF REPAYMENT OF NOTE PAYABLE
Date Interest expense principal repayment Carrying value of note
December 31 2020 1089102
2021 1089102*9%= 98019.18 280000-98019.18= 181980.82 1089102-181980.82= 907121.18
2022 907121.18*9%= 81640.91 280000-81640.91= 198359.09 907121.18-198359.09= 708762.09
2023 708762.09*9%= 63788.59= 280000-63788.59= 216211.41 708762.09-216211.41=492550.68
2024 492550.68*9%= 44329.56 280000-44329.56= 235670.44 492550.68-235670.44= 256880.24
2025 256880.24*9%= 23119.76 280000-23119.76= 256880.24 256880.24-256880.24 = 0

3)

Date Account title Debit credit
31December 2020 Machinery 1889102
cash 800000
Note payable 1089102
31December 2021 Note payable 181980.82
Interest expense 98019.18
cash 280000
31December 2022 Note payable 198359.09
Interest expense 81640.91
cash 280000
31December 2023 Note payable 216211.41
Interest expense 63788.59
cash 280000
31December 2024 Note payable 235670.44
Interest expense 44329.56
cash 280000
31December 2025 Note payable 256880.24
Interest expense 23119.76
cash 280000